# Generating Completions for Fragmented Broca’s Aphasic Sentences Using Large Language Models

### v1

@@ Abstract
Broca’s aphasia is a type of aphasia characterized by non-fluent, effortful and fragmented speech production with relatively good comprehension. Since traditional aphasia treatment methods are often time-consuming, labour-intensive, and do not reflect real-world conversations, applying natural language processing based approaches such as Large Language Models (LLMs) could potentially contribute to improving existing treatment approaches. To address this issue, we explore the use of sequence-to-sequence LLMs for completing fragmented Broca’s aphasic sentences. We first generate synthetic Broca’s aphasic data using a rule-based system designed to mirror the linguistic characteristics of Broca’s aphasic speech. Using this synthetic data, we then fine-tune four pre-trained LLMs on the task of completing fragmented sentences. We evaluate our fine-tuned models on both synthetic and authentic Broca’s aphasic data. We demonstrate LLMs’ capability for reconstructing fragmented sentences, with the models showing improved performance with longer input utterances. Our result highlights the LLMs’ potential in advancing communication aids for individuals with Broca’s aphasia and possibly other clinical populations.
{textblock}
1.6(0,3.2) This work has been submitted to the IEEE for possible publication.
Copyright may be transferred without notice, after which this version may no longer be accessible.
@@
Aphasia, assistive technology, fragmented speech, generative AI, large language models, non-fluent speech, synthetic data generation.
@@ 1 Introduction and Background
Aphasia is an acquired language disorder after a focal brain injury. Typically, aphasia has long-term consequences on the communication abilities of the affected individuals, and thus, negatively impacts their quality of life [1, 2] together with their relatives’ quality of life [3]. Within the spectrum of aphasia, Broca’s aphasia, the target group for this study, is a sub-type of aphasia described as having non-fluent, effortful, telegraphic speech production, but with relatively preserved comprehension. Given the significant language fragmentation in Broca’s aphasic speech, diverse treatment approaches have been proposed [4]. While studies have demonstrated the effectiveness of these treatments, mixed findings regarding long-term maintenance of gains post-treatment [5, 6] have led to the emergence of alternative communication aids. These alternative approaches, although beneficial, present some limitations such as inefficiency, lack of personalization, and poor generalization to real-world conversations [3]. In spontaneous interactions with spouses and caregivers, individuals with Broca’s aphasia engage in more fluid communication, highlighting the need for interventions that better address the dynamics of natural language use in everyday settings. The goal of the current study is to investigate the potential of reconstructing fragmented Broca’s aphasic language into full utterances by using Large Language Models (LLMs) fine-tuned on synthetic Broca’s aphasic data [^1] The code and data are available online: https://github.com/sijbrenvv/Completions_for_Broca-s_aphasia.
@@ 1.1 Linguistic characteristics of Broca’s aphasic language
Despite the difficulties in production, comprehension abilities in Broca’s aphasia remain relatively unaffected. This discrepancy suggests that linguistic representations may be preserved in people with aphasia (PWA), but access to these representations is impaired [11, 15, 12]. In summary, individuals with Broca’s aphasia produce highly fragmented speech, with one- to two-word utterances in more severe cases.
@@ 1.2 AI applications in aphasiology
Previous research has primarily focused on using NLP techniques for the automatic detection of aphasia (with potential application in diagnostic purposes), while largely overlooking the use of these techniques for developing assistive technologies to enhance communication for PWA, their spouses and caregivers. A recent study [22] reported an innovative approach that could potentially contribute to the advancement of assistive technology by using LLMs for reconstructing fragmented Broca’s aphasic sentences. The goal of that study was twofold. First, LLMs require large amounts of training data, while large datasets of aphasic speech do not exist. Therefore, the study [22] proposes a system that takes as an input a corpus of written texts from healthy speakers (C4 dataset [23]), and generates synthetic aphasic utterances by modifying the sentences in the corpus so that their linguistic characteristics match those reported for Broca’s aphasic speech. Second, the authors then use this synthetic dataset to fine-tune a pre-trained LLM, specifically the T5 model [23], on a sentence completion task. Although the study [22] showed promising results, it presents a few methodological problems. First, the written text in the C4 corpus (used to generate synthetic aphasic utterances) does not accurately represent daily spoken language, potentially limiting the real-life applicability of the resulting system. Second, even though the study [22] aims to demonstrate the potential of synthetic datasets for developing communication aids for PWA, the lack of testing on real aphasic utterances makes it unclear to what extent the presented approach is applicable to actual clinical scenarios. Lastly, the study’s use of the C4 dataset for generating synthetic aphasic sentences and evaluating the fine-tuned LLM causes a potentially serious issue of data leakage, since the C4 dataset was included in the original training data for the T5 LLM. A data leakage could lead to artificially inflated performance metrics and even compromise the validity of the model’s evaluation and its future application on real aphasic speech.
@@
The present study aims to address these problems by answering the following research question: To what extent can a pretrained sequence-to-sequence LLM, fine-tuned on synthetic Broca’s aphasic data, effectively reconstruct authentic fragmented sentences produced by individuals with Broca’s aphasia into complete, grammatically correct sentences? To address this research question, we first generate synthetic fragmented Broca’s aphasic sentences based on spoken language sentences produced by neurotypical individuals. This is done using a rule-based system adapted from the above-mentioned study [22]. Second, we use our synthetic data to fine-tune a series of pretrained LLMs on a sentence completion task, and then evaluate these LLMs on synthetic and authentic aphasic sentences.
@@ 2.1 Synthetic data generation
Discard 50 % regular $\rightarrow\O$
@@
Ideally, our synthetic sentences should be hard to discriminate from authentic aphasic sentences, and both of them should be easy to discriminate from sentences produced by neurotypical individuals. To test whether this is the case, we use a series of standard machine learning binary classification tasks. We first prepare three datasets containing (1) neurotypical and authentic aphasic sentences, (2) neurotypical and synthetic aphasic sentences, and (3) authentic and synthetic aphasic sentences. In each dataset, the two types of sentences are mixed in equal proportions. We then split each dataset into a training and test portions using 80–20 ratio. Using each dataset, we train a Naive Bayes classifier to discriminate between the two types of sentences, and present the evaluation results in Section 3.1 below.
@@ 2.2.2 Fine-tuning setup
We use the HuggingFace transformer library [30] to load and quantise each model to 4-bit precision with a batch size of 8. For efficiency, we employ HuggingFace’s Parameter-Efficient Fine-Tuning (PEFT) library [31] that incorporates Quantized Low-Rank Adaptation (QLoRA) [32] (see more details in Appendix .1).
Using these techniques, we fine-tune each model. T5-base is fine-tuned for a maximum of 5 epochs with early stopping after 2 epochs with no improvement, using a learning rate of $10^{-4}$ with half precision (FP16). The other models are fine-tuned for $1$ epoch only due to overfitting. Appendix .2 provides technical details about fine-tuning each model.
To evaluate our sentence completion models, we use two different methods. First, we compare the models’ completions of (unseen) synthetic sentences to their original counterparts (i.e., the same sentences before they are processed by our synthetic data generation system). This comparison is done using two measures: ChrF between the two sentences (mentioned above) and cosine similarity between the embeddings of the two sentences (using embeddings of [33]). While ChrF measures the degree of surface-level (i.e., character) alignment between the two sentences, cosine similarity measures how close the meanings of the two sentences are. Second, we use our models to generate completions for authentic sentences produced by PWA. In this case, we have no ground-truth data to compare our models’ completions to, and we limit our evaluation to qualitative analyses.
@@ 2.3.3 Neurotypical individuals’ data from SBCSAE
The Santa Barbara Corpus of Spoken American English (SBCSAE) [35] is a compilation of naturally occurring spoken interactions containing approximately 250k word tokens. SBCSAE is comparable to AphasiaBank in that it consists of transcriptions of spoken language. Finally, the corpus is very diverse and representative, as the data was collected from individuals of varied regional origin, age, ethnic and social background, occupation, and gender. We use utterances from SBCSAE as input to our synthetic data generation system.
The sentences in both AphasiaBank and SBCSAE include linguistic annotations, special characters, and other types of noise. We preprocess the data sources to extract only the actual word tokens and remove all annotations (i.e., pauses, actions, text within angle and square brackets), unicode errors and special characters, instances of stuttering and repeated words, etc. A full list of preprocessing routines with examples can be found in Appendix .3.
@@ 3.1 Evaluating synthetic data generation system
Table 3: Accuracy of a Naive Bayes classifier in binary classification tasks.
Scenario Sent. pairs Accuracy
Authentic aphasic vs. neurotypical 9,861 0.80
Synthetic aphasic vs. neurotypical 7,998 0.81
Authentic aphasic vs. synthetic aphasic 7,998 0.83
To evaluate our synthetic data generation system, we look at the accuracy of a Naive Bayes classifier in three different binary classification tasks: (1) authentic aphasic vs. neurotypical, (2) synthetic aphasic vs. neurotypical, and (3) authentic vs. synthetic aphasic sentences. Ideally, we expect a low accuracy score on the task of discriminating synthetic vs. authentic aphasic sentences, and high accuracy scores on the tasks of discriminating either of them vs. neurotypical sentences. From Table 3, we indeed observe high accuracy scores for (either authentic or synthetic) aphasic vs. neurotypical sentences, indicating that a linear classifier easily distinguishes between aphasic and neurotypical utterances. This result meets our expectations. The comparable accuracy scores on the two tasks (0.80 vs. 0.81) suggest that both authentic and synthetic aphasic utterances differ from neurotypical utterances to a similar extent.
Meanwhile, the classifier also easily differentiates between authentic and synthetic Broca’s aphasic utterances, which implies that they are also linearly separable to a high degree. While this is not what we expected, this result does not necessarily mean that the synthetic data generated by our system is unsuitable for fine-tuning LLMs. The mechanics of LLM fine-tuning go far beyond simple linear classification, and whether our synthetic data helps LLM fine-tuning is an empirical question. Since our generated data appears satisfactory upon manual inspection, we proceed with using the synthetic Broca’s aphasic data generated by our system to fine-tune the four LLMs.
@@ 3.2.1 Synthetic data
In this setting, we compare the fragmented (synthetic) sentences completed by our four LLMs to the original sentences.
@@
Model ChrF Cos. sim. ChrF Cos. sim.
T5-base 56.39$\pm$0.48 0.88$\pm$0.00 56.56$\pm$0.48 0.88$\pm$0.00
Flan-T5-base 49.53$\pm$0.42 0.87$\pm$0.00 50.15$\pm$0.42 0.87$\pm$0.00
T5-large 53.95$\pm$0.46 0.88$\pm$0.00 54.94$\pm$0.47 0.88$\pm$0.00
Flan-T5-XL 55.71$\pm$0.47 0.88$\pm$0.00 56.19$\pm$0.48 0.88$\pm$0.00
@@
A visual inspection of the generated completions (see examples in Table 5) reveals that the models, especially T5-base and Flan-T5-XL, show great capacity in addressing some of the main deficits observed in Broca’s aphasic utterances, such as word omission. While T5-base achieves the highest performance scores, it occasionally generates semantically incorrect completions, which highlights limitations in its ability to generate contextually appropriate sentences. Flan-T5-XL, on the other hand, performs consistently well, often opting for simpler but semantically plausible completions. This tendency reveals a trade-off between complexity and fluency, and suggests that reconstructing more morphosyntactically complex utterances is a challenging task for our models. More examples of generated completions for the synthetic test set are presented in Appendix .4, Table 12.
@@ 3.2.2 Authentic data
Table 6 (top half) presents examples of completions generated by the two LLMs. Although the models exhibit some capability of handling Broca’s aphasic speech, important limitations remain. In particular, both models sometimes only minimally change the authentic utterance or even simply reproduce it. Also, they occasionally add negations or expand contractions, which can change the sentence meaning to the opposite.
While there are instances of grammatical completions, the models often produce outputs that do not correct the grammatical errors present in the authentic Broca’s aphasic utterances. This indicates a gap in the models’ ability to fully interpret and rectify the Broca’s aphasic speech patterns. Both models occasionally insert common conversational fillers or affirmations such as “Yeah”, “Okay” and “I do not know”, which do not meaningfully contribute to completing the utterance but reflect an attempt to provide a plausible conversational continuation.
There are some utterances so fragmented that completing them meaningfully would be challenging even for humans. In these cases, the models either do not change the utterances at all or provide completions that do not improve clarity. For example, “Peanut, b@you,.” by T5-base remains largely unreadable.
Very short utterances might be difficult for our models to complete not only because they are often cryptic, but also because our models were not fine-tuned on this task: recall from Section 2.1 above that our synthetic utterances had to contain at least three words. Table 6 (bottom half) presents examples of completions for longer and more complex utterances (three words or more, see Appendix .5, Table 14, for more examples).
For such longer utterances, the models demonstrate a better grasp of context, producing more complete and semantically coherent responses, with less redundancy and fewer nonsensical completions. Even for uncommon phrases, the models are better at generating plausible continuations. For instance, “Oh is it hid feet” was completed to “Oh is it hid under the feet” by T5-base, which, despite a minor grammatical error, shows an effort to logically complete the utterance.

### v2

# Generating Completions for Broca’s Aphasic Sentences Using Large Language Models

@@ Abstract
Broca’s aphasia is a type of aphasia characterized by non-fluent, effortful and agrammatic speech production with relatively good comprehension. Since traditional aphasia treatment methods are often time-consuming, labour-intensive, and do not reflect real-world conversations, applying natural language processing based approaches such as Large Language Models (LLMs) could potentially contribute to improving existing treatment approaches. To address this issue, we explore the use of sequence-to-sequence LLMs for completing Broca’s aphasic sentences. We first generate synthetic Broca’s aphasic data using a rule-based system designed to mirror the linguistic characteristics of Broca’s aphasic speech. Using this synthetic data (without authentic aphasic samples), we then fine-tune four pre-trained LLMs on the task of completing agrammatic sentences. We evaluate our fine-tuned models on both synthetic and authentic Broca’s aphasic data. We demonstrate LLMs’ capability for reconstructing agrammatic sentences, with the models showing improved performance with longer input utterances. Our result highlights the LLMs’ potential in advancing communication aids for individuals with Broca’s aphasia and possibly other clinical populations.
@@
Aphasia, assistive technology, agrammatic speech, generative AI, large language models, non-fluent speech, synthetic data generation.
@@ 1 Introduction and Background
Aphasia is an acquired language disorder after a focal brain injury. Typically, aphasia has long-term consequences on the communication abilities of the affected individuals, and thus, negatively impacts their quality of life [1, 2] together with their relatives’ quality of life [3]. Within the spectrum of aphasia, Broca’s aphasia, the target group for this study, is a sub-type of aphasia described as having non-fluent, effortful, telegraphic speech production, but with relatively preserved comprehension. Given the significant language fragmentation in Broca’s aphasic speech, diverse treatment approaches have been proposed [4]. While studies have demonstrated the effectiveness of these treatments, mixed findings regarding long-term maintenance of gains post-treatment [5, 6] have led to the emergence of alternative communication aids. These alternative approaches, although beneficial, present some limitations such as inefficiency, lack of personalization, and poor generalization to real-world conversations [3]. Existing communication aids for people with aphasia (PWA) include low-tech options like picture or spelling boards and high-tech devices with pre-defined messages (such as making an order for coffee), but these tools often limit expression to basic ideas or messages and can be challenging for PWA to navigate [7, 8, 9, 10, 11, 3]. Furthermore, these alternative communication aids frequently require specific skills, such as keyboard use or navigating pictograms, which can pose significant challenges for many PWA, thereby restricting their ability to engage in daily conversations [3]. Navigating pictogram sets or relying on pre-defined texts often does not fully capture the contextual dependencies inherent in real-world conversations. Consequently, developing interventions that can replicate the fluid, context-dependent language patterns observed in spontaneous interactions with familiar partners, such as spouses and caregivers, is of considerable importance. The goal of the current study is to investigate the potential of reconstructing agrammatic language into full utterances by using Large Language Models (LLMs) fine-tuned on synthetic Broca’s aphasic data [^1] The code and data are available online: https://github.com/sijbrenvv/Completions_for_Broca-s_aphasia.
@@ 1.1 Linguistic characteristics of Broca’s aphasic language
Despite the difficulties in production, comprehension abilities in Broca’s aphasia remain relatively unaffected. This discrepancy suggests that linguistic representations may be preserved in PWA, but access to these representations is impaired [16, 20, 17]. In summary, individuals with Broca’s aphasia produce highly agrammatic speech, with one- to two-word utterances in more severe cases.
@@ 1.2 AI applications in aphasiology
Previous research has primarily focused on using NLP techniques for the automatic detection of aphasia (with potential application in diagnostic purposes), while largely overlooking the use of these techniques for developing assistive technologies to enhance communication for PWA, their spouses and caregivers. Recent scoping reviews underscore significant gaps in the application of AI within aphasiology, particularly in its therapeutic and assistive uses [27, 28, 29]. One review [27] reported that 7 out of their 77 reviewed studies explored applications of AI in therapeutic or assistive applications, while most research focused on automated assessment and diagnosis of aphasia. The review highlighted a predominant reliance on supervised machine learning for classification tasks, with limited use of advanced AI techniques for personalized rehabilitation or self-management. Similarly, [29] did not identify any studies integrating AI into augmentative and alternative communication (AAC) devices specifically tailored for aphasia rehabilitation. Together, these findings emphasize a disproportionate focus on diagnostic applications over functional communication support, revealing a critical need for AI-driven innovations that address real-world conversational challenges faced by PWA and their caregivers.
@@
At the same time, [30] recently reported an innovative approach that could potentially contribute to the advancement of assistive technology by using LLMs for reconstructing agrammatic aphasic sentences. The goal of that study was twofold. First, LLMs require large amounts of training data, while large datasets of aphasic speech do not exist. Therefore, the study [30] proposes a system that takes as an input a corpus of written texts from healthy speakers (C4 dataset [31]), and generates synthetic aphasic utterances by modifying the sentences in the corpus so that their linguistic characteristics match those reported for Broca’s aphasic speech. Second, the authors then use this synthetic dataset to fine-tune a pre-trained LLM, specifically the T5 model [31], on a sentence completion task. Although the study [30] showed promising results, it presents a few methodological problems. First, the written text in the C4 corpus (used to generate synthetic aphasic utterances) does not accurately represent daily spoken language, potentially limiting the real-life applicability of the resulting system. Second, even though the study [30] aims to demonstrate the potential of synthetic datasets for developing communication aids for PWA, the lack of testing on real aphasic utterances makes it unclear to what extent the presented approach is applicable to actual clinical scenarios. Lastly, the study’s use of the C4 dataset for generating synthetic aphasic sentences and evaluating the fine-tuned LLM causes a potentially serious issue of data leakage, since the C4 dataset was included in the original training data for the T5 LLM. A data leakage could lead to artificially inflated performance metrics and even compromise the validity of the model’s evaluation and its future application on real aphasic speech.
@@
The present study aims to address these problems by answering the following research question: To what extent can a pretrained sequence-to-sequence LLM, fine-tuned on synthetic Broca’s aphasic data, effectively reconstruct authentic agrammatic sentences produced by individuals with Broca’s aphasia into complete, grammatically correct sentences? To address this research question, we first generate synthetic Broca’s aphasic sentences based on spoken language sentences produced by neurotypical individuals. This is done using a rule-based system adapted from the above-mentioned study [30]. Second, we use our synthetic data to fine-tune a series of pretrained LLMs on a sentence completion task, and then evaluate these LLMs on synthetic and authentic aphasic sentences.
@@ 2.1 Synthetic data generation
Discard 50 % regular $\rightarrow\varnothing$
@@
Ideally, our synthetic sentences should be similar to authentic aphasic sentences and different from sentences produced by neurotypical individuals. There are many possible ways to compare sentence characteristics, but what matters in our setup is that the patterns described above hold in the context of LLMs. This is why we estimate the average acceptability (or predictability) of a sample of sentences in each data set using two language models commonly used for this purpose, GPT-2 [36] and RoBERTa [37]. Predictability is estimated through the standard measure called surprisal, or negative log-probability [38, 39], which measures how easy it is on average for a language model to predict words in a sentence given their context. We apply this approach to sentences produced by neurotypical individuals, PWA, and our synthetic data generation system. We only consider sentences that are longer than 3 words and downsample each data set to 4,448 sentences (the size of the smallest, authentic aphasic, data set after filtering out short sentences). We expect that the language produced by neurotypical individuals should be highly predictable and, therefore, associated with lower surprisal, while both authentic and synthetic authentic data should be more irregular, and, therefore, less predictable and associated with higher surprisal.
@@ 2.2.2 Fine-tuning setup
We use the HuggingFace transformer library [42] to load and quantise each model to 4-bit precision with a batch size of 8. For efficiency, we employ HuggingFace’s Parameter-Efficient Fine-Tuning (PEFT) library [43] that incorporates Quantized Low-Rank Adaptation (QLoRA) [44] (see more details in Supplementary Materials). Using these techniques, we fine-tune each model. T5-base is fine-tuned for a maximum of 5 epochs with early stopping after 2 epochs with no improvement, using a learning rate of $10^{-4}$ with half precision (FP16). The other models are fine-tuned for $1$ epoch only due to overfitting. Supplementary Materials provide technical details about fine-tuning each model.
To evaluate our sentence completion models, we use two different methods. First, we compare the models’ completions of (unseen) synthetic sentences to their original counterparts (i.e., the same sentences before they are processed by our synthetic data generation system). This comparison is done using three measures: ChrF and RougeL [45] between the two sentences (mentioned above), and cosine similarity between the embeddings of the two sentences (using embeddings of [46]). While ChrF and RougeL measure the degree of surface-level (i.e., character and word, respectively) alignment between the two sentences, cosine similarity measures how close the meanings of the two sentences are. Although BLEU [47] is a popular evaluation metric in sentence comparison tasks, we found it to be unsuitable for our sentence completion task. The explanation with examples can be found in Supplementary Materials. Second, we use our models to generate completions for authentic sentences produced by PWA. In this case, we have no ground-truth data to compare our models’ completions to, and we limit our evaluation to qualitative analyses.
@@ 2.3.3 Neurotypical individuals’ data from SBCSAE
The Santa Barbara Corpus of Spoken American English (SBCSAE) [49] is a compilation of naturally occurring spoken interactions containing approximately 250k word tokens. SBCSAE is comparable to AphasiaBank in that it consists of transcriptions of spoken language. Finally, the corpus is very diverse and representative, as the data was collected from individuals of varied regional origin, age, ethnic and social background, occupation, and gender. We use utterances from SBCSAE to provide input to our synthetic data generation system, as well as to evaluate this system.
The sentences in both AphasiaBank and SBCSAE include linguistic annotations, special characters, and other types of noise. We preprocess the data sources to extract only the actual word tokens and remove all annotations (i.e., pauses, actions, text within angle and square brackets), unicode errors and special characters, instances of stuttering and repeated words, etc. A full list of preprocessing routines with examples can be found in Supplementary Materials.
@@ 3.1 Evaluating synthetic data generation system
Table 3: Average sentence predictability (surprisal) assigned to each data type by two pretrained LLMs. While the synthetic aphasic sentences are not directly extracted from any specific source, they are generated based on the SBCSAE data.
Data type Source GPT-2 RoBERTa
Neurotypical SBCSAE 4.76 2.96
Neurotypical AphasiaBank 5.43 3.50
Authentic aphasic AphasiaBank 6.48 5.08
Synthetic aphasic – (SBCSAE) 6.39 5.51
To evaluate our synthetic data generation system, we consider the average sentence predictability (surprisal) values shown in Table 3. Note that GPT-2 and RoBERTa have very different architectures and are trained on different data sets, therefore, the surprisal values across the two models (columns) are not directly comparable, and we only compare the values within each model (column). We first notice that the two data sets that contain sentences from neurotypical individuals (rows 1–2) are associated with lower surprisal values compared to the authentic aphasic data (row 3), and this result holds for the values generated by both language models, the pattern we expected. Second, our synthetic data (row 4) yields surprisal values that are comparable to the values for the authentic aphasic data (row 3) and higher than those for the neurotypical data (rows 1–2). This is an important result that suggests that our synthetic generation system produces sentences where words are harder for LLMs to predict from context – similar to authentic aphasic sentences and in contrast to neurotypical language. Upon manual inspection, our generated data also appears satisfactory, therefore, we proceed with using the synthetic Broca’s aphasic data generated by our system to fine-tune the four LLMs.
@@ 3.2.1 Synthetic data
In this setting, we compare the agrammatic (synthetic) sentences completed by our four LLMs to the original sentences.
@@
Model ChrF RougeL Cos. sim. ChrF RougeL Cos. sim.
[…]
@@
A visual inspection of the generated completions (see examples in Table 5) reveals that the models, especially T5-base and Flan-T5-XL, show great capacity in addressing some of the main deficits observed in Broca’s aphasic utterances, such as word omission. While T5-base achieves the highest performance scores, it occasionally generates semantically incorrect completions, which highlights limitations in its ability to generate contextually appropriate sentences. Flan-T5-XL, on the other hand, performs consistently well, often opting for simpler but semantically plausible completions. This tendency reveals a trade-off between complexity and fluency, and suggests that reconstructing more morphosyntactically complex utterances is a challenging task for our models. More examples of generated completions for the synthetic test set are presented in Supplementary Materials.
@@ 3.2.2 Authentic data
Table 6 (top half) presents examples of completions generated by the two LLMs. Although the models exhibit some capability of handling Broca’s aphasic speech, important limitations remain. In particular, both models sometimes only minimally change the authentic utterance or even simply reproduce it. Also, they occasionally add negations (2.99% of the time for FLAN-T5-XL), which can change the sentence meaning to the opposite, or expand contractions wrongly (e.g., “He’s” to “He is” instead of “He has”). The occasional addition of negation reflects the general difficulty of this phenomenon for natural language processing models [50, 51].
While there are instances of grammatical completions, the models often produce outputs that do not correct the grammatical errors present in the authentic Broca’s aphasic utterances. This indicates a gap in the models’ ability to fully interpret and rectify the Broca’s aphasic speech patterns. Both models occasionally insert common conversational fillers or affirmations such as “Yeah”, “Okay” and “I do not know”, which do not meaningfully contribute to completing the utterance but reflect an attempt to provide a plausible conversational continuation. There are some utterances so agrammatic that completing them meaningfully would be challenging even for humans. In these cases, the models either do not change the utterances at all or provide completions that do not improve clarity. For example, “Peanut, b@you,.” by T5-base remains largely unreadable.
To estimate how well FLAN-T5-XL performs in the target task, we have manually annotated 10% of our test data, namely 455 sentences. We did the annotation without considering their broader context, to simulate the task that the model was faced with. We found that 196 sentences did not require any completion by the model, while 259 did. From the sentences that needed to be completed, 66 would not be possible to complete, given the lack of the broader context, while 193 could in principle be completed. Out of these, the model has successfully completed 80 sentences, simply reproduced 57 of them, and produced errors in 56 sentences (42 cases of incomplete/minimal change, 12 unnecessary additions, and 2 cases of added negation). On top of that, the model introduced unnecessary changes into the 13 (out of the 196) sentences that did not require completion.
Very short utterances might be difficult for our models to complete not only because they are often cryptic, but also because our models were not fine-tuned on this task: recall from Section 2.1 above that our synthetic utterances had to contain at least three words. Table 6 (bottom half) presents examples of completions for longer and more complex utterances (three words or more, see Supplementary Materials for more examples). For such longer utterances, the models demonstrate a better grasp of context, producing more complete and semantically coherent responses, with less redundancy and fewer nonsensical completions. Even for uncommon phrases, the models are better at generating plausible continuations. For instance, “Oh is it hid feet” was completed to “Oh is it hid under the feet” by T5-base, which, despite a minor grammatical error, shows an effort to logically complete the utterance.